# Using iterated local alignment to aggregate trajectory data into a traffic flow map

### v5

@@ 1 Introduction
Figure 1: Trajectory data in Hannover, Germany. (a) Empirical trajectories (green circles). Klusreide road segment (solid yellow circle). (b) Zoom of trajectories and map matched routes (blue lines), with road network (solid black lines), near Klusreide. (d) Zoom of 2 m $\times$ 2 m rasterised flow aggregation. Colour scale is traffic flow in grid cells.
@@
If we zoom in on these trajectory points around Klusriede (solid yellow circle), a residential street near the A2 autobahn, in Figure 1(b), then we observe that the points deviate from the road network (solid black lines). These noisy trajectories are unsuitable for direct aggregation into a flow map, so the usual first step is to apply a map matching procedure (Quddus et al., 2007, Chao et al., 2020). The map matched routes are the blue lines in Figure 1(b). These follow the road network more closely. Map matching is a highly difficult problem and map matching with exact alignment to the underlying road network in a reasonable time frame is usually not feasible. The approach for fast computations taken by the Valhalla routing engine introduces some approximations which do not lead to exact alignment of empirical trajectories with the road network or to each other (Sghaier, 2019). Nonetheless the misalignment is reduced to less than 5 m, which is, for many purposes, including most notably navigation, indeed a suitable result.
@@
For these inexactly aligned trajectories, applying the state-of-the-art traffic flow aggregation API (Morgan and Lovelace, 2021) leads to an inaccurate flow map. The authors resolve this by rasterising the flow map to a raster grid: and within each raster grid cell, the flow is the sum of the flows of all trajectories which intersect the grid cell. This rasterised flow map is displayed in Figure 1(c). The colour scale for the flow ranges from 100 (violet) to 300 (pink) to 600 (orange). The 2 m $\times$ 2 m raster grid obscures the road network at this scale. The resolution of a rasterised aggregation flow map is fixed by the resolution of the raster grid. If we require a higher resolution, then we must recalculate the flow map on a raster grid with a higher resolution.
@@
Our goal is to produce a traffic flow map from map matched routes, which can be utilised at any scale without recalculation, from the city level to the individual road level. In our proposed local alignment approach, we infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it. To accomplish this, we introduce several novel algorithms which employ a mix of advanced statistical and geospatial methods. These include “node snapping” where the nearby boundary points of road segments are combined via statistical clustering, and “line blending” where road segments, which are near each other but do not share overlapping sub-segments, are aligned to maximise their overlapping sub-segments. We combine these novel algorithms into an iterative workflow which outputs a locally aligned flow map. This flow map is composed of road segments rather than raster grid cells, so its resolution is not limited by the raster grid resolution.
@@ 2 Local alignment of route segments for flow aggregation
We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that different road segments are accessible to/from each other at this node point. We set $\mathcal{N}$ to be the OpenStreetMap (OSM) road network. Each path in the network is composed of a connected sequence of edges, and we refer to this as a “linestring” geometry, following the Open Geospatial Consortium (OGC) terminology (OGC, 2010). We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$. This is a “multipoint” geometry (OGC, 2010). The points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$.

### v6

@@ 1 Introduction
Figure 1: Trajectory data in Hannover, Germany. (a) Empirical trajectories (green circles), Loebensteinstraße (solid cyan circle). (b) Zoom of trajectories and map matched routes (blue lines). (c) Zoom of ideal flow aggregation. (d) Zoom of existing flow aggregation. (e) Zoom of 2 m $\times$ 2 m rasterised flow aggregation. Colour scale/label is traffic flow in road segments/grid cells.
@@
If we zoom in on these trajectory points around Loebensteinstraße (solid cyan circle), formerly known as Hindenburgstraße, is a residential street near the A2 autobahn, then we observe that in Figure 1(b) the points deviate from the road network. These noisy trajectories are unsuitable for direct aggregation into a flow map, so the usual first step is to apply a map matching procedure (Quddus et al., 2007, Chao et al., 2020). The map matched routes are the blue lines in Figure 1(b). These follow the road network more closely. Map matching is a highly difficult problem and map matching with exact alignment to the underlying road network in a reasonable time frame is usually not feasible. The approach for fast computations taken by the Valhalla routing engine introduces some approximations which do not lead to exact alignment of empirical trajectories with the road network or to each other (Sghaier, 2019). Nonetheless the misalignment is reduced to less than 5 m, which is, for many purposes, including most notably navigation, indeed a suitable result.
@@
The ideal flow aggregation from the empirical trajectories is in Figure 1(c). The colour scale for the flow ranges from 0 (violet) to 150 (pink) to 300 (yellow). Applying the state-of-the-art traffic flow aggregation API (Morgan and Lovelace, 2021) to the inexactly aligned trajectories leads to an inaccurate flow map in Figure 1(d). The authors resolve this by rasterising the flow map to a raster grid: within each raster grid cell, the flow is the sum of the flows of all trajectories which intersect the grid cell. This rasterised flow map is displayed in Figure 1(e). The 2 m $\times$ 2 m raster grid obscures the road network at this scale. The resolution of a rasterised aggregation flow map is fixed by the resolution of the raster grid. If we require a higher resolution, then we must recalculate the flow map on a raster grid with a higher resolution.
@@
Our goal is to produce a traffic flow map from empirical trajectories which is aligned to the road network and which can be utilised at any scale without recalculation, from the city level to the individual road level. In our proposed local alignment approach, we infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it. To accomplish this, we introduce a novel algorithm “line blending” where road segments, which are near each other but do not share overlapping sub-segments, are aligned to maximise their overlapping sub-segments. Line blending is a “fuzzy” aggregation method which calculates accurate flow aggregation of noisy trajectories. This is in contrast to the current state-of-the-art flow aggregation methods which are accurate only for noise-free trajectories. We combine these novel algorithms into an iterative workflow which outputs a locally aligned flow map. This flow map is composed of road segments rather than raster grid cells, so its resolution is not limited by the raster grid resolution.
@@ 2 Local alignment of route segments for flow aggregation
We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges $\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that different road segments are accessible to/from each other at this node point. We set $\mathcal{N}$ to be the OpenStreetMap (OSM) road network. Each path in the network is composed of a connected sequence of edges, and we refer to this as a “linestring” geometry. We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$. This is a “multipoint” geometry. The points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$.